Let w be a negative real number, z a 6th root of w.

a) Show that z(k) = p^1/6 [cos ( pi + 2 * k* pi / 6) + i sin ( pi + 12 * k * pi / 6)], k = v, 1,2,3,4,5 is a formula for the 6th roots of w.

b) Hence determine 6th roots of - 729.

c) Given z = cosA + i sin A and u + iv = (1+z) (1+ z^2). Prove that v = u tan(3A/2), u^2 + v^2 = 16 cos ^2(A/2) cos^2A
1
Expert's answer
2020-05-04T19:19:46-0400

Let us assume that "w=-p", where "p" is a positive real number. Next we use the trigonometric form of complex number

"p= \rho(\cos\varphi + i\sin\varphi)."

For positive real numbers "\varphi = 0" and "\rho=p" , so

"p = p(\cos 0 + i\sin 0)."

Therefore

"w=-p=-p(\cos 0 + i\sin 0) = p(\cos\pi + i\sin\pi),"

because "\cos\pi = -\cos 0" and "\sin\pi = \sin 0."

a) Let us assume that

"z = \rho(\cos\theta+i\sin\theta)"

According to de Moivre's formula,

"z^6 = \rho^6(\cos(6\theta)+i\sin(6\theta))".

Since "z" is a 6th root of "w",

"z^6=w"

and

"\rho^6(\cos(6\theta)+i\sin(6\theta)) = p(\cos\pi +i\sin\pi)."

Therefore we get "\rho^6=p" , "\cos(6\theta)=\cos\pi" , "\sin(6\theta)=\sin\pi." Then we obtain

"\rho=p^{1/6},

6\theta=\pi +2\pi k, k \in \mathbb{Z}."

Finally, we get

"\rho=p^{1/6},

\theta(k)=\dfrac{\pi +2\pi k}{6}, k \in \mathbb{Z}."

We can see that we will obtain the same "\sin\theta, \cos\theta" for "k\equiv_6 m" , so we consider only "k=0,1,\ldots 5."

Next we obtain

"z(k)=p^{1/6}(\cos(\frac{\pi+2\pi k}{6}) + i\sin(\frac{\pi+2\pi k}{6})), k = 0,\,\ldots, 5."

b) Here "w=-729" . In trigonometric form

"w=729(\cos\pi + i\sin\pi),"

according to the formula in a) 6th roots are

"z(k)= 729^{1/6}(\cos(\frac{\pi+2\pi k}{6}) + i\sin(\frac{\pi+2\pi k}{6})), k = 0,\,\ldots, 5."

Therefore,

"z(0)=3(\cos(\frac{\pi}{6}) + i\sin(\frac{\pi}{6})) = 3(\frac{\sqrt{3}}{2}+\frac12i),"

"z(1)= 3(\cos(\frac{\pi+2\pi}{6}) + i\sin(\frac{\pi+2\pi}{6})) = 3(0+i)= 3i,"

"z(2)=3(\cos(\frac{\pi+4\pi}{6}) + i\sin(\frac{\pi+4\pi}{6})) = 3(-\frac{\sqrt3}{2} +\frac12i),"

"z(3)=3(\cos(\frac{\pi+6\pi}{6}) + i\sin(\frac{\pi+6\pi}{6})) = 3(-\frac{\sqrt3}{2} -\frac12i),"

"z(4)=3(\cos(\frac{\pi+8\pi}{6}) + i\sin(\frac{\pi+8\pi}{6})) = 3(0-i) = -3i,"

"z(5)=3(\cos(\frac{\pi+10\pi}{6}) + i\sin(\frac{\pi+10\pi}{6})) = 3(\frac{\sqrt3}{2} -\frac12i)."

c) Let us consider "w_1=1+z = (1+\cos A) + i\sin A." The modulus of "w_1" is

"\rho_1=\sqrt{(1+\cos A)^2+(\sin A)^2} = \sqrt{1+2\cos A + \cos^2A + \sin^2A} = \sqrt{2+2\cos A}."

Let us consider "w_2=1+z^2 = (1+ \cos2A) + i\sin 2A." The modulus of "w_2" is

"\rho_2=\sqrt{(1+\cos 2A)^2+(\sin2 A)^2} = \sqrt{1+2\cos 2A + \cos^22A + \sin^22A} = \sqrt{2+2\cos 2A}."

If we use the trigonometric form, the modulus of a product is equal to the product of moduli.

The sum "u^2+v^2" is the squared modulus of "w_1\cdot w_2," therefore

"u^2+v^2 = \rho_1^2\rho_2^2 = (2+2\cos A)\cdot(2+2\cos 2A) = 16(\dfrac{1+\cos A}{2})\cdot(\dfrac{1+\cos 2A}{2}) = 16\cos^2\frac{A}{2}\cos^2A."

Next we'll simplify the formula

"u+iv = (1+z)(1+z^2) = 1+z+z^2+z^3."

According to de Moivre's formula,

"1+z+z^2+z^3 = 1+ (\cos A + i\sin A) + (\cos 2A+i\sin 2A) + (\cos 3A+i\sin3A) = (1+\cos A+ \cos 2A + \cos 3A) + i\cdot(\sin A+ \sin2A+\sin3A)."

Therefore,

"u = 1+\cos A+ \cos 2A + \cos 3A, v=\sin A+ \sin2A+\sin3A."

Using several trigonometric identities, we obtain

"u = 1 + \cos A+ \cos 2A + \cos 3A = 1 + \cos A + 2\cos^2A-1+ 4\cos^3A-3\cos A= 4\cos^3A+2\cos^2A-2\cos A = 2\cos A \cdot(\cos A+1)\cdot(2\cos A -1)."

Next we obtain

"v=\sin A+ \sin2A+\sin3A = \sin A + 2\sin A\cos A + 3\sin A(1-\sin^2A) - \sin^3A = 2 \cos A\cdot\sin A\cdot( 2\cos A +1)."

Let us consider "\tan(3A/2)":

"\tan\dfrac{3A}{2} = \dfrac{\sin3A}{1+\cos3A} = \dfrac{(2\cos A+1)\cdot\sin A}{(\cos A+1)\cdot(2\cos A-1)}."

We should prove that "v = u \tan(\frac{3A}{2})" . Let us substitute the expressions obtained above:

"u\tan(\frac{3A}{2}) = 2\cos A\cdot(\cos A +1) \cdot(2\cos A - 1) \cdot \dfrac{(2\cos A+1)\cdot\sin A}{(\cos A+1)\cdot(2\cos A-1)}= 2\cos A\cdot (2\cos A+1)\cdot\sin A = v."